Stealth aircraft are designed to not reflect radar, whose wavelength is typically 2 cm, by using a antireflecting coating. Ignoring any change in wavelength in the coating, estimate its thickness.
Answer #1

Use the following formula to find the thickness of the coating,

\(2 t=\frac{\lambda}{2}\)

Here, \(t\) is the thickness and \(\lambda\) is the wavelength and \(\frac{\lambda}{2}\) is the path difference.

The thickness of the coating is,

\(t=\frac{\lambda}{2}\)

\(=\frac{2 \mathrm{~cm}}{4}\)

\(=0.5 \mathrm{~cm}\)